# TTF-VLA: Temporal Token Fusion via Pixel-Attention Integration for Vision-Language-Action Models

### v1

@@ 1 Introduction
Vision-Language-Action (VLA) models have emerged as a transformative paradigm in robotic manipulation, seamlessly integrating visual perception, natural language understanding, and action generation within unified neural architectures. Building upon the success of large-scale vision-language transformers, recent VLA systems (Brohan et al. 2022, 2023; Octo Model Team et al. 2024; Kim et al. 2024; Kim, Finn, and Liang 2025; Black et al. 2024, 2025) have demonstrated unprecedented capabilities in executing complex manipulation instructions across diverse environments. These models fundamentally reshape robotic control by treating action prediction as a multimodal sequence generation task, where visual observations and natural language instructions are jointly processed to produce discrete action tokens that guide continuous robot behavior.
@@
  • Comprehensive experimental validation: 4.0 percentage points average improvement on LIBERO (Liu et al. 2023), cross-environment generalization on SimplerEnv (Li et al. 2024), and 8.7% relative improvement on real robot tasks, with model-agnostic applicability across OpenVLA (Kim et al. 2024) and VLA-Cache (Xu et al. 2025) architectures.
@@ 2 Related Work
Vision-Language-Action Models Vision-Language-Action (VLA) models have emerged as a unified framework for robotic manipulation, integrating visual perception, language understanding, and action prediction. Early works like RT-1 and RT-2 tokenized actions with visual and linguistic inputs for end-to-end policy learning (Brohan et al. 2022, 2023). Recent advancements include Octo and OpenVLA for open-source implementations (Octo Model Team et al. 2024; Kim et al. 2024; Kim, Finn, and Liang 2025), and Pi-0/Pi-0.5 for flow-based architectures (Black et al. 2024, 2025). However, current models exhibit limitations in handling temporal dynamics and visual artifacts (Chi et al. 2023; Wen et al. 2024; Collaboration et al. 2023; Huang et al. 2024), often processing frames independently while ignoring temporal coherence.
Token Processing Techniques Efficient token processing has become crucial for transformer deployment. In vision transformers, attention-guided methods such as DynamicViT, AdaViT, and EViT progressively prune redundant tokens (Rao et al. 2021; Meng et al. 2022; Liang et al. 2022), while ToMe optimizes efficiency through strategic token consolidation (Bolya et al. 2023). For vision-language models, recent approaches like FastV and SparseVLM focus on visual token sparsification using text-guided selection strategies (Chen et al. 2024; Zhang et al. 2024). VLA-Cache specifically targets robotic scenarios through KV-cache reuse mechanisms (Xu et al. 2025). However, these methods primarily address spatial redundancy within individual frames. Unlike these spatial compression approaches, our work targets temporal redundancy across sequential frames, proposing a dual-dimension strategy that leverages historical information to enhance VLA inference quality.
@@ 3 Methodology
Figure 1: Overall Framework of Temporal Token Fusion for VLA Models. The framework illustrates the end-to-end process, where the Vision Encoder extracts tokens from current (Observationt) and previous (Observationt-1) frames. These are then processed by the Patch Selection module and TTF module for patch selection and token fusion. The fused tokens are subsequently fed into the LLM Backbone, combined with language instruction, to generate 7D robotic actions via the Action De-Tokenizer.
@@ 3.3 Dual-Dimension Detection
The pixel-level dimension captures fine-grained spatial changes through efficient grayscale-based analysis (Fasola and Veloso 2006), contributing to the identification of important patches as part of the dual-dimension detection in Figure 2(a). Our approach offers several advantages over token-space similarity metrics commonly used in caching systems.
@@
This conversion is motivated by robotic manipulation scenarios where meaningful changes (object movements, shadow variations, lighting shifts) are primarily reflected in luminance rather than chromatic information.(Yang, Tan, and Ahuja 2012; Smeulders et al. 2014; Secci and Ceccarelli 2023)
@@
The attention dimension identifies semantically important patches using transformer attention patterns. Building upon token selection methods in vision-language models (Chen et al. 2024; Zhang et al. 2024) and attention-guided approaches in vision transformers (Rao et al. 2021; Meng et al. 2022; Liang et al. 2022; Bolya et al. 2023; Xu et al. 2025), our approach introduces dual attention sources specifically adapted for temporal fusion in robotic manipulation tasks.
@@ 4.1 Experimental Setup
  • Real Robot Tasks: Physical validation using a Franka Research 3 robot across three manipulation tasks spanning different complexities: single-object pick-and-place (put the garlic on the plate), multi-object sequential manipulation (put the pepper and corn on the plate), and contact-rich manipulation (close the drawer). We collect 80 demonstration episodes per task using Gello teleoperation at 5 Hz, then fine-tune OpenVLA-7B for 20,000 steps using batch size 8 across 8 A100 GPUs. Evaluation deploys the fine-tuned models on a single A100 at 5 Hz with 20 episodes per task.
@@ 4.2 Main Results
Table 1 presents our core experimental results across four task suites using two representative VLA architectures, demonstrating the effectiveness and model-agnostic applicability of our temporal token fusion approach. Our analysis reveals several key insights: Model-agnostic effectiveness: Both OpenVLA and VLA-Cache demonstrate consistent improvements (4.0 and 2.7 percentage points average respectively), validating the generalizability of our approach across different VLA architectures. Task-specific patterns: Long-horizon tasks benefit most significantly from temporal fusion (+11.5% relative improvement for OpenVLA), suggesting that extended manipulation sequences particularly benefit from temporal context integration, while object manipulation shows strong absolute gains (+6.0 percentage points for OpenVLA). Fusion efficiency: The fusion rates (42.8% average for OpenVLA) indicate substantial feature reuse while maintaining performance gains, demonstrating that our dual-dimension detection successfully identifies stable regions for temporal integration without compromising inference quality. Temporal progression analysis: Figure 3 illustrates representative failure-to-success cases where baseline methods fail but TTF succeeds, highlighting the critical role of temporal coherence in successful robotic manipulation task completion.
@@
Implicit Query Reuse and Noise Robustness: A particularly revealing insight emerges from VLA-Cache+TTF results. VLA-Cache (Xu et al. 2025) accelerates inference by reusing Key-Value matrices ($\mathbf{K}_{t-1}^{(l)},\mathbf{V}_{t-1}^{(l)}$) for static visual patches while always recomputing Query matrices ($\mathbf{Q}_{t}^{(l)}$) to maintain contextual sensitivity. However, when TTF is applied to VLA-Cache, an important mechanism emerges: TTF’s token-level fusion ($\tilde{\mathbf{t}}_{t}^{(i)}=\mathbf{t}_{t-1}^{(i)}$ for selected patches) means that the corresponding Query matrix portions are implicitly reused since $\mathbf{Q}_{t}^{(l)}=\mathbf{W}_{q}^{(l)}\cdot\tilde{\mathbf{T}}_{t}$. Thus, VLA-Cache+TTF effectively reuses all three attention matrices (K, V, and Q) for static patches, going beyond VLA-Cache’s original KV-only reuse strategy. Contrary to conventional wisdom that Query reuse degrades performance due to contextual sensitivity, our results show significant improvements, particularly for long-horizon tasks (+3.0 points: 55.0%→58.0%). This performance improvement demonstrates that our dual-dimension detection provides stabilized contextual representations that enhance robustness against visual observation noise—including lighting fluctuations, motion blur, and sensor artifacts common in robotic manipulation. The selective Query reuse validates TTF’s core hypothesis: temporal coherence in stable regions enhances rather than compromises inference quality. This finding reveals a promising “free lunch” future direction worth exploring: direct KQV matrix reuse for static patches could achieve computational acceleration while simultaneously improving task success rates.
@@
Configuration Strategy: Parameter selection demonstrates TTF’s robustness. OpenVLA employs consistent parameters: keyframe interval K=3, attention top-k=70, pixel threshold=0.03, and text-to-vision attention. VLA-Cache requires minimal adaptation with task-adaptive fusion rates (30% for Object/Spatial/Long, 50% for Goal tasks). Since VLA-Cache uses patch importance scoring, we directly utilize its scores for patch selection, avoiding redundant computation. Neither model requires task-specific hyperparameter tuning, with comprehensive parameter exploration detailed in Supplementary Material.
@@ 5 Conclusion
We present Temporal Token Fusion (TTF), a training-free approach that addresses the fundamental limitation of current VLA models in leveraging temporal coherence during robotic manipulation. Our dual-dimension detection framework combines pixel-level spatial dynamics with attention-based semantic relevance assessment to intelligently integrate historical and current visual representations. Comprehensive evaluation demonstrates consistent improvements across LIBERO (4.0 percentage points average), SimplerEnv (cross-environment generalization), and real robot tasks (8.7% relative improvement), with model-agnostic applicability across OpenVLA and VLA-Cache architectures. Beyond performance gains, our work reveals the unexpected benefits of selective Query matrix reuse in attention mechanisms, suggesting promising directions for direct KQV matrix strategies that achieve computational acceleration while improving task success rates. The approach is particularly effective for long-horizon tasks, where temporal context integration provides the most substantial benefits, demonstrating enhanced robustness against visual noise common in real-world robotic environments.

### v3

@@ 1 Introduction
Vision-Language-Action (VLA) models have emerged as a transformative paradigm in robotic manipulation, seamlessly integrating visual perception, natural language understanding, and action generation within unified neural architectures. Building upon the success of large-scale vision-language transformers, recent VLA systems (brohan2022rt1; brohan2023rt2; pertsch2024octo; kim2024openvla; kim2025oft; black2024pi0; black2025pi05) have demonstrated unprecedented capabilities in executing complex manipulation instructions across diverse environments. These models fundamentally reshape robotic control by treating action prediction as a multimodal sequence generation task, where visual observations and natural language instructions are jointly processed to produce discrete action tokens that guide continuous robot behavior.
@@
  • Comprehensive experimental validation: 4.0 percentage points average improvement on LIBERO (liu2023libero), cross-environment generalization on SimplerEnv (li24simpler), and 8.7% relative improvement on real robot tasks, with model-agnostic applicability across OpenVLA (kim2024openvla) and VLA-Cache (xu2025vla) architectures.
@@ 2 Related Work
Vision-Language-Action Models Vision-Language-Action (VLA) models have emerged as a unified framework for robotic manipulation, integrating visual perception, language understanding, and action prediction. Early works like RT-1 and RT-2 tokenized actions with visual and linguistic inputs for end-to-end policy learning (brohan2022rt1; brohan2023rt2). Recent advancements include Octo and OpenVLA for open-source implementations (pertsch2024octo; kim2024openvla; kim2025oft), and Pi-0/Pi-0.5 for flow-based architectures (black2024pi0; black2025pi05). However, current models exhibit limitations in handling temporal dynamics and visual artifacts (chi2023diffusion; wen2024diffusion; openx2023; huang2024a3vlm), often processing frames independently while ignoring temporal coherence.
Token Processing Techniques Efficient token processing has become crucial for transformer deployment. In vision transformers, attention-guided methods such as DynamicViT, AdaViT, and EViT progressively prune redundant tokens (dynamicvit2021; adavit2022; evit2022), while ToMe optimizes efficiency through strategic token consolidation (tome2023). For vision-language models, recent approaches like FastV and SparseVLM focus on visual token sparsification using text-guided selection strategies (chen2024image; zhang2024sparsevlm). VLA-Cache specifically targets robotic scenarios through KV-cache reuse mechanisms (xu2025vla). However, these methods primarily address spatial redundancy within individual frames. Unlike these spatial compression approaches, our work targets temporal redundancy across sequential frames, proposing a dual-dimension strategy that leverages historical information to enhance VLA inference quality.
@@ 3 Methodology
Figure 1: Overall Framework of Temporal Token Fusion for VLA Models. The framework illustrates the end-to-end process, where the Vision Encoder extracts tokens from current (Observationt) and previous (Observationt-1) frames. These are then processed by the Patch Selection module and TTF module for patch selection and token fusion. The fused tokens are subsequently fed into the LLM Backbone, combined with language instruction, to generate 7-DoF robotic actions via the Action Detokenizer.
@@ 3.3 Dual-Dimension Detection
The pixel-level dimension captures fine-grained spatial changes through efficient grayscale-based analysis (fasola2006real), contributing to the identification of important patches as part of the dual-dimension detection in Figure 2(a). Our approach offers several advantages over token-space similarity metrics commonly used in caching systems.
@@
This conversion is motivated by robotic manipulation scenarios where meaningful changes (object movements, shadow variations, lighting shifts) are primarily reflected in luminance rather than chromatic information.(yang2012shadow; smeulders2014visual; secci2023rgb)
@@
The attention dimension identifies semantically important patches using transformer attention patterns. Building upon token selection methods in vision-language models (chen2024image; zhang2024sparsevlm) and attention-guided approaches in vision transformers (dynamicvit2021; adavit2022; evit2022; tome2023; xu2025vla), our approach introduces dual attention sources specifically adapted for temporal fusion in robotic manipulation tasks.
@@ 4.1 Experimental Setup
  • Real Robot Tasks: Physical validation using a Franka Research 3 robot across three manipulation tasks spanning different complexities: single-object pick-and-place (put the garlic on the plate), multi-object sequential manipulation (put the pepper and corn on the plate), and contact-rich manipulation (close the drawer). We collect 80 demonstration episodes per task using Gello teleoperation at 5 Hz, and fine-tune OpenVLA-7B for 20,000 steps with a batch size of 8. During evaluation, the fine-tuned models are deployed at 5 Hz and tested with 20 episodes for each task.
@@ 4.2 Main Results
Table 1 presents our core experimental results across four task suites using two representative VLA architectures, demonstrating the effectiveness and model-agnostic applicability of our temporal token fusion approach. Our analysis reveals several key insights: Model-agnostic effectiveness: Both OpenVLA and VLA-Cache demonstrate consistent improvements (4.0 and 2.7 percentage points average respectively), validating the generalizability of our approach across different VLA architectures. Task-specific patterns: Long-horizon tasks benefit most significantly from temporal fusion (+11.5% relative improvement for OpenVLA), suggesting that extended manipulation sequences particularly benefit from temporal context integration, while object manipulation shows strong absolute gains (+6.0 percentage points for OpenVLA). Fusion efficiency: The fusion rates (42.8% average for OpenVLA) indicate substantial feature reuse while maintaining performance gains, demonstrating that our dual-dimension detection successfully identifies stable regions for temporal integration without compromising inference quality. Temporal progression analysis: Figure 3 illustrates representative failure-to-success cases where baseline methods fail but TTF succeeds, highlighting the critical role of temporal coherence in successful robotic manipulation task completion. TTF introduces less than 2% additional runtime overhead, confirming its efficiency and suitability for real-time robotic inference.
@@
Implicit Query Reuse and Noise Robustness: A particularly revealing insight emerges from VLA-Cache+TTF results. VLA-Cache (xu2025vla) accelerates inference by reusing Key-Value matrices ($\mathbf{K}_{t-1}^{(l)},\mathbf{V}_{t-1}^{(l)}$) for static visual patches while always recomputing Query matrices ($\mathbf{Q}_{t}^{(l)}$) to maintain contextual sensitivity. However, when TTF is applied to VLA-Cache, an important mechanism emerges: TTF’s token-level fusion ($\tilde{\mathbf{t}}_{t}^{(i)}=\mathbf{t}_{t-1}^{(i)}$ for selected patches) means that the corresponding Query matrix portions are approximately and implicitly reused since $\mathbf{Q}_{t}^{(l)}=\mathbf{W}_{q}^{(l)}\cdot\tilde{\mathbf{T}}_{t}$. Thus, VLA-Cache+TTF effectively nearly reuses all three attention matrices (K, V, and Q) for static patches, going beyond VLA-Cache’s original KV-only reuse strategy. Contrary to conventional wisdom that Query reuse degrades performance due to contextual sensitivity, our results show significant improvements, particularly for long-horizon tasks (+3.0 points: 55.0%→58.0%). This performance improvement demonstrates that our dual-dimension detection provides stabilized contextual representations that enhance robustness against visual observation noise—including lighting fluctuations, motion blur, and sensor artifacts common in robotic manipulation. The selective Query reuse validates TTF’s core hypothesis: temporal coherence in stable regions enhances rather than compromises inference quality. This finding reveals a promising “free lunch” future direction worth exploring: direct KQV matrix reuse for static patches could achieve computational acceleration while simultaneously improving task success rates.
@@
Configuration Strategy: Parameter selection demonstrates TTF’s robustness. OpenVLA employs consistent parameters: keyframe interval K=3, attention top-k=70, pixel threshold=0.03, and text-to-vision attention. VLA-Cache requires minimal adaptation with task-adaptive fusion rates (30% for Object/Spatial/Long, 50% for Goal tasks). Since VLA-Cache uses patch importance scoring, we directly utilize its scores for patch selection, avoiding redundant computation. Neither model requires task-specific hyperparameter tuning.
@@ 5 Conclusion
We present Temporal Token Fusion (TTF), a training-free approach that addresses the fundamental limitation of current VLA models in leveraging temporal coherence during robotic manipulation. Our dual-dimension detection framework combines pixel-level spatial dynamics with attention-based semantic relevance assessment to intelligently integrate historical and current visual representations. Comprehensive evaluation demonstrates consistent improvements across LIBERO (4.0 percentage points average), SimplerEnv (cross-environment generalization), and real robot tasks (8.7% relative improvement), with model-agnostic applicability across OpenVLA and VLA-Cache architectures. Beyond performance gains, our work reveals the unexpected benefits of selective Query matrix reuse in attention mechanisms, suggesting promising directions for direct KQV matrix strategies that achieve computational acceleration while improving task success rates.